What is the concentration of NaOH solution for which 11.25 mL is needed to titrate 1.53 g of benzoic acid?



1
Expert's answer
2020-12-11T02:48:20-0500

Mass of Benzoic acid= 1.53g

Molar mass of C6H5COOH= (12x6)+5+16+16+1= 110gmol-1

moles of Benzoic acid= mass/molar mass= 1.53/110=0.014mol

The balanced chemical equation of reaction is

C6H5COOH + NaOH ---> C6H5COONa + H2O

1mol of Benzoic acid reacts with 1 mole of NaOH

Therefore, 0.014mol of Benzoic acid will react with 0.014mol of NaOH.

Moles of NaOH= 0.014mol

Volume of NaOH= 11.25ml = 11.25/1000 =0.01125L

Concentration of NaOH (Molarity)= moles of NaOH/ Litres of solution

= 0.014mol/0.01125L

= 1.24moldm-3

Therefore, the concentration of NaOH is 1.24moldm-3